OpenStudy (anonymous):

A suit was on sale for $142.50, marked down 25% from its original price. What was the original price of the suit?

OpenStudy (anonymous):

let n be the original price. so, (75/100)n=142.50. try solve this equation to find value of n.

OpenStudy (anonymous):

or u can try 1.25 X 142.50

OpenStudy (anonymous):

u got the answer WOMH?

OpenStudy (anonymous):

178.125

OpenStudy (anonymous):

the answer should be 190.

OpenStudy (anonymous):

yep, i messed up

OpenStudy (anonymous):

since there are 25% discount. this means that the current price is 75% of the original price. or in fraction, it is 3/4 of the original price. let n denote the original price, then u wil have an equation like this. (3/4)n= 142.50 therefore, n=142.50*(4/3) = 190

OpenStudy (anonymous):

thanks so much!